OpenStudy (anonymous):

4(x - 6) = -2(x - 3) plz show ur work!

OpenStudy (anonymous):

have you tried the problem yourself...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think that u have to use the distributive problem

OpenStudy (anonymous):

like 4times x making 4x-24=2x-12

OpenStudy (anonymous):

just multiply it all out then simplify

OpenStudy (anonymous):

well ya there you go

OpenStudy (anonymous):

and then 4x-2x making 2x - 12

OpenStudy (anonymous):

and than 12 divided by 2 making x 6

OpenStudy (anonymous):

simplify then you get your answer

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

x= 6 yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wait isnt it negative 6

OpenStudy (anonymous):

wait ur wrong

OpenStudy (anonymous):

and then 4x-2x making 2x - 12 i was just using this :P

OpenStudy (anonymous):

but lets see

OpenStudy (anonymous):

i multiplied wrong

OpenStudy (anonymous):

the answer is 9

OpenStudy (anonymous):

4(x - 6) = -2(x - 3) 4x -24 = -2x + 6 <------ here we go :D

OpenStudy (anonymous):

simplify that

OpenStudy (anonymous):

ya the answer should not be 9

OpenStudy (anonymous):

unless i copied the problem wrong

OpenStudy (anonymous):

2x-18 2x/2 and 18/2 is 9

OpenStudy (anonymous):

4(x - 6) = -2(x - 3) 4x -24 = -2x + 6 4x - 24 + (24) = - 2x + 6 + ( 24 ) additive inverse of 24 24 is canceled on the rt side , added to right (simplified) 4x = -2x + 30 4x + (-2x) = -2x + 30 + (2x ) additive inverse of 2x -2x is cancelled on lt side, added to lt (simplified) 6x = 30 do you understand why it is not 9

OpenStudy (anonymous):

i think you can do the rest on your own

OpenStudy (anonymous):

@lubavasilyuk

OpenStudy (anonymous):

its up to you if you want to believe it

OpenStudy (anonymous):

x is 5